# Remarks on nonperturbative perturbations

### v2

@@ 1 Introduction
Here, we focus on the extremal Reissner-Nordström black hole in $d$-dimensions, which has a near-horizon $\mathrm{AdS}_{2}\times S^{d-2}$ geometry. In this context, Ref. Horowitz:2023xyl considered linearized tensor perturbations in the near-horizon region, whose functional form is entirely determined by symmetry considerations, decomposing into harmonics of $\mathrm{AdS}_{2}$ and $S^{d-2}$. This reduces the equations of motion for the linearized perturbations to a system of algebraic equations. Of particular note, the perturbations scale as $r^{\gamma}$ for some scaling dimension $\gamma$, where $r$ is the radial coordinate of the $\mathrm{AdS}_{2}$. As long as $\gamma$ is a non-negative integer, such perturbations are analytic as we approach the horizon at $r=0$. However, for non-integer $\gamma$, the metric is only $C^{\lfloor\gamma\rfloor}$. The equations of motion effectively reduce to a fourth-order polynomial equation for $\gamma$, with two of the roots being physically meaningful solutions. This was done for the case of the general four-derivative action
@@
In particular, regardless of the sign of $d_{0}$, the higher-derivative corrections generically lead to non-integer $\gamma$, and hence non-analytic behavior on the horizon. This non-analyticity was interpreted in Horowitz:2023xyl as a breakdown of effective field theory, so it is natural to ask if these instabilities persist if we treat the higher-derivative corrections nonperturbatively.
At the very least, such a resolution works for the thermodynamics. There are divergences in the thermodynamic quantities associated with higher-derivative extremal black holes, but it was noticed by Ref. Wu:2024iiz that this is due to the perturbative nature of the corrections. In particular, the authors looked at the nonperturbative static black hole solutions Wiltshire:1985us of Einstein-Maxwell-Gauss-Bonnet gravity in five dimensions and found that the perturbative corrections to the temperature, entropy, and electric potential all scaled as $\alpha(\mu^{2}-q^{2})^{-1}$ relative to the leading order, where $\mu$ and $q$ are integration constants roughly corresponding to mass and charge, $\alpha$ is the Gauss-Bonnet coupling, and $\mu=q+\tfrac{\alpha}{2}$ is the extremality condition. However, the full nonperturbative solution admits an extremal limit, which can then be expanded perturbatively in $\alpha$ without issue. This suggests that the problem arises from not being careful about $\mu^{2}-q^{2}$ becoming $\mathcal{O}(\alpha)$. Indeed, when $\mu^{2}-q^{2}\propto\alpha$, the thermodynamic quantities receive an infinite series of $\mathcal{O}(\alpha^{0})$ corrections.
This motivates us to ask if nonperturbative higher-derivative corrections might fix the issue of diverging tidal forces in asymptotically flat space. In this short note, we consider the case of Einstein-Maxwell-Gauss-Bonnet gravity in $d$-dimensions and find the nonperturbative scaling dimensions. The expressions themselves are not particularly enlightening, so we leave them to the main text and appendices. However, we find that the nonperturbative corrections are generically non-analytic and hence do not fix the instability. This suggests that the instability is not merely an order of limits issue but a genuine breakdown of effective field theory. We also perform the calculation of nonperturbative scaling dimensions for the asymptotically AdS case for completeness, although it clearly cannot resolve the instability as it is already present at the two-derivative level.
@@ 2 Flat Gauss-Bonnet black holes
where $\check{g}$ and $\check{A}$ denote the perturbed solution while $g$ and $A$ denote the original static solution (5). We will work to linear order in the perturbations. Due to the symmetry, such perturbations decompose into harmonics of $\mathrm{AdS}_{2}$ and $S^{d-2}$,
@@
As noted in Horowitz:2022mly , $\gamma_{--}$ and $\gamma_{-+}$ are always negative and hence blow up at the horizon, and therefore may be removed by appropriate boundary conditions on the horizon.
@@
In particular, we see that the $\ell=2$ mode has a scaling dimension
@@
which results in tidal force singularities on the horizon. Thus, the instability persists nonperturbatively in $\alpha$.
@@ 4 Lovelock black holes
Here we focus on the case of third-order Lovelock gravity ($N=3$) in asymptotically flat space ($\alpha_{0}=0$). Moreover, we specialize to the case $d=7$ as this is the lowest dimension for which $\mathcal{X}_{6}$ is non-trivial. Explicitly,
@@
The $\mathcal{O}(\alpha)$ term is always parametrically smaller than the $\mathcal{O}(\alpha^{0})$ term, regardless of $\ell$. However, the $\mathcal{O}(\alpha^{2})$ term dominates for large $\ell$, and hence we see that the perturbative expansion breaks down at $\ell\sim\alpha_{2}^{-1}$, up to $\mathcal{O}(1)$ numerical factors. A similar breakdown was observed in the Eikonal limit for Kerr black holes in Cano:2024bhh . [^3] Note that the Eikonal limit also involves taking the azimuthal parameter $m\to\infty$ with $\ell/m$ fixed, and so differs from the limit we take. In any case, we cannot directly compare to four dimensions since all the higher-derivative Lovelock invariants become trivial.
@@ 5 Discussion
We have computed the nonperturbative (in higher-derivative couplings) scaling dimensions of the (linearized) perturbations of near-horizon Reissner-Nordström black holes in Einstein-Maxwell-Gauss-Bonnet gravity and third-order Lovelock gravity. In particular, the horizon instabilities of Horowitz:2022mly ; Horowitz:2023xyl ; Horowitz:2024dch remain in this analysis. However, to our knowledge, this is the first computation of nonperturbative (in higher-derivative couplings) scaling dimensions for black hole perturbations, which is interesting in its own right.

### v3

@@ 1 Introduction
Here, we focus on the extremal Reissner-Nordström black hole in $d$-dimensions, which has a near-horizon $\mathrm{AdS}_{2}\times S^{d-2}$ geometry. In this context, Ref. Horowitz:2023xyl considered linearized tensor perturbations in the near-horizon region, whose functional form is entirely determined by symmetry considerations, decomposing into harmonics of $\mathrm{AdS}_{2}$ and $S^{d-2}$. This reduces the equations of motion for the linearized perturbations to a system of algebraic equations. Of particular note, the perturbations scale as $r^{\gamma}$ for some scaling dimension $\gamma$, where $r$ is the radial coordinate of the $\mathrm{AdS}_{2}$. As long as $\gamma$ is a non-negative integer, such perturbations are analytic as we approach the horizon at $r=0$. However, for non-integer $\gamma$, the metric is only $C^{\lfloor\gamma\rfloor}$, and, for non-zero $\gamma<1$, the metric suffers from a tidal force singularity at the horizon. The equations of motion effectively reduce to a fourth-order polynomial equation for $\gamma$, with two of the roots being physically meaningful solutions. This was done for the case of the general four-derivative action
@@
In particular, regardless of the sign of $d_{0}$, the higher-derivative corrections generically lead to non-integer $\gamma$, and hence non-analytic behavior on the horizon. For the particular case of $d=\ell+3$, the two-derivative $\gamma_{+-}$ is zero and the sign is fully determined by the higher-derivative corrections. For $d_{0}>0$, this leads to a much stronger singular behavior on the horizon.
This non-analyticity was interpreted in Horowitz:2023xyl as a breakdown of effective field theory, so it is natural to ask if these instabilities persist if we treat the higher-derivative corrections nonperturbatively. Of course, we expect there to be an infinite series of higher-derivative corrections, which is computationally intractable to deal with. However, for certain theories, namely those that are quasitopological, we can treat a finite number of higher-derivative corrections non-perturbatively. In particular, this gives us a toy model with which we may explore whether or not the instabilities are resolved. In this short note, we consider the case of Einstein-Maxwell-Gauss-Bonnet gravity in $d$-dimensions and find the nonperturbative scaling dimensions. The expressions themselves are not particularly enlightening, so we leave them to the main text and appendices. However, we find that the nonperturbative corrections are generically non-analytic and hence do not fix the instability. This suggests that the instability is a genuine breakdown of effective field theory. We also perform the calculation of nonperturbative scaling dimensions for the asymptotically AdS case for completeness, although it clearly cannot resolve the instability, as it is already present at the two-derivative level.
@@ 2 Flat Gauss-Bonnet black holes
Due to the quasi-topological nature of the theory, it does not suffer from the infamous Ostrogradsky instability that plagues higher-derivative theories, even if $\alpha$ is treated nonperturbatively. This also has the consequence that the equations of motion are second-order in derivatives
@@
where $\check{g}$ and $\check{A}$ denote the perturbed solution while $g$ and $A$ denote the original static solution (5). We will work to linear order in static, scalar-derived perturbations. Due to the symmetry, such perturbations decompose into harmonics of $\mathrm{AdS}_{2}$ and $S^{d-2}$,
@@
As noted in Horowitz:2022mly , we are always free to choose boundary conditions on the horizon to remove two of the four solutions. $\gamma_{--}$ and $\gamma_{-+}$ are always negative and hence blow up at the horizon, and therefore we will choose to remove these, in keeping with the conventions of Horowitz:2022mly . It should be noted that for $d>5$, $\gamma_{+-}$ is already negative at the two-derivative level when $\ell<d-3$. For such solutions, the linearized ansatz breaks down near the horizon.
@@
For example, we see that the $\ell=2$ mode has a scaling dimension
@@
which generically results in tidal force singularities on the horizon. For large enough $\ell$, there is no tidal force singularity but the higher-derivative corrections still do not resolve the non-analyticity. Thus, the instability persists nonperturbatively in $\alpha$.
It is curious to observe that the $\ell=2$ scaling dimension is negative when $\alpha<0$, although this does not seem to be a physical choice of sign. Moreover, when $\gamma<0$, the metric blows up at the horizon, and we should not trust the linearized expansion. In the extreme case that $\alpha$ becomes negative and large enough, the scaling dimension becomes complex. Additionally, since we are working nonperturbatively in $\alpha$, we are free to take $\alpha$ large. In particular, $\alpha=n(n+1)r_{+}^{2}/6$ corresponds to $\gamma_{+-}=n$ (for $\ell=2$ and $n\geq 0$), so we can technically resolve the instability for large $\alpha$. Of course, this only resolves one instability since $\alpha$ can’t depend on $\ell$.
@@ 4 Lovelock black holes
Here we focus on the case of third-order Lovelock gravity ($N=3$) in asymptotically flat space ($\alpha_{0}=0$), and we specialize to the case $d=7$, which is the lowest dimension for which $\mathcal{X}_{6}$ is non-trivial. As before, we expect the physical choice of sign to be $\alpha_{1}>0$, but, to our knowledge, the positivity bounds on $\alpha_{2}$ are not known. Explicitly,
@@
The $\mathcal{O}(\alpha)$ term is always parametrically smaller than the $\mathcal{O}(\alpha^{0})$ term, regardless of $\ell$. However, the $\mathcal{O}(\alpha^{2})$ term dominates for large $\ell$, and hence we see that the perturbative expansion breaks down at $\ell\sim\alpha_{2}^{-1}$, up to $\mathcal{O}(1)$ numerical factors. A similar breakdown was observed in the Eikonal limit for Kerr black holes in Cano:2024bhh . [^4] Note that the Eikonal limit also involves taking the azimuthal parameter $m\to\infty$ with $\ell/m$ fixed, and so differs from the limit we take. In any case, we cannot directly compare to four dimensions since all the higher-derivative Lovelock invariants become trivial. It should be noted that (46) is not valid for $\ell=2$, which, if treated carefully, has the expansion
$\gamma_{+-}=-\frac{1}{2}+\frac{3}{2r_{+}}\sqrt{\frac{21\alpha_{2}}{2}}-\sqrt{42}\quantity(1197+832\frac{\alpha_{3}}{\alpha_{2}^{2}})\frac{\alpha_{2}^{3/2}}{r_{+}^{3}}+\mathcal{O}(\alpha^{3}).$ (47)
Nevertheless, $\gamma_{+-}$ is negative at the two-derivative level for $\ell=2$, so the linearized perturbation result should not be trusted.
@@ 5 Discussion
We have computed the nonperturbative (in higher-derivative couplings) scaling dimensions of the (linearized) perturbations of near-horizon Reissner-Nordström black holes in Einstein-Maxwell-Gauss-Bonnet gravity and third-order Lovelock gravity. In particular, the horizon instabilities of Horowitz:2022mly ; Horowitz:2023xyl ; Horowitz:2024dch remain in this analysis. This does not necessarily rule out the possibility of there being some particular choice of (an infinite series of) higher-derivative corrections that manages to make the instability disappear, but it does seem to suggest that such resolution would have to be fine-tuned rather than generically arising from a nonperturbative treatment of the corrections. However, to our knowledge, this is the first computation of nonperturbative (in higher-derivative couplings) scaling dimensions for black hole perturbations, which is interesting in its own right.